OpenStudy (anonymous):

if f(x) = 3/x+2 - square root x-3, complete the following statement F(19)=?

hartnn (hartnn):

did u put x=19 in f(x) what do u get ?

OpenStudy (anonymous):

19=3/21- squareroot 19-3

OpenStudy (anonymous):

19=3/21-4?

OpenStudy (anonymous):

is the answer 3?

hartnn (hartnn):

3/21 is 1/7 actually, so, f(19) = 1/7 -4 or if you simplify = -27/7